# Improving Data Efficiency for LLM Reinforcement Fine-tuning Through Difficulty-targeted Online Data Selection and Rollout Replay

### v1

@@ 5.1 Experimental Setup
We perform GRPO training on three model scales: Qwen2.5-Math-1.5B, Qwen2.5-3B, and Qwen2.5-Math-7B [46]. We adopt four open-source datasets for training: MATH [13], DeepScaleR-40K [27], Open-Reasoner-Zero-57K (ORZ) [15] and DeepMath-103K dataset [12]. For MATH, we include all level 3–5 questions. For the other three datasets, we sample 8K to 10K subsets to construct the training pools. These datasets span diverse mathematical domains and difficulty levels. In total, we experiment with six LLM-training dataset combinations to assess the effectiveness of our framework.
@@
In practice, we observe that off-the-shelf pretrained embedding models struggle to capture fine-grained similarity between math questions. To address this, we freeze the Qwen2.5-Math-1.5B-Instruct backbone [46] and train a 3-layer MLP adapter with a calibration head, using binary cross-entropy loss. We fix the reference set size to 256. Additional implementation details are provided in Appendix D.1 and ablation results on key components are presented in Appendix F.1.
@@
We employ the verl framework [40] to perform GRPO training. We use a batch size of 512, generate 8 rollouts per prompt, and train for 60 steps in all experiments. The maximum rollout length is set to 3072 tokens for the Qwen2.5-Math series models (due to max position embedding limits) and 4096 tokens for Qwen2.5-3B. For reward computation, we use a simple rule-based function based solely on answer correctness, without incorporating any format-related signals. The 1.5B and 3B models are trained on 8 L40S GPUs, while the 7B model is trained on 8 A100 GPUs. For DOTS, data selection is performed every 2 steps. For RR, we choose the fresh rollout ratio $\delta$ as $0.5$ and buffer capacity $C\in\{256,512\}$. All RL training hyperparameters are detailed in Appendix E.2.
@@
Figure 3: Average accuracy curves of our method and original GRPO under various LLM–dataset combinations. The curves show average performance aggregated over four benchmarks with exponential smoothing for visualization. Although both methods are trained for the same number of steps (60), our curve is shorter in duration because RR reduces the wall-clock time per step. Our method consistently outperforms the original GRPO throughout training and reduces the time required to match the original GRPO’s final accuracy after 60 training steps by an average of 38.8%.
@@ 5.2 Main Results
The total training costs can be decomposed into two components: the number of steps required to reach a target performance and the average wall-clock time per step. Each training step consists of processing a fixed-size batch, primarily including rollout generation and policy update. To ensure a fair comparison, each set of experiments is run on the same type and number of GPU devices.
Our method reaches the same performance as original GRPO with fewer steps.
Tab. 1 reports the number of training steps required by DOTS+RR to match the final performance of the original GRPO at 60 steps. Across all LLM–dataset combinations, our method consistently reaches the same performance with substantially fewer steps, achieving reductions ranging from 13.33% to 60.00%. These results demonstrate that DOTS significantly accelerates convergence by prioritizing informative training samples.
@@
In our experiments, rollout generation accounts for approximately 47%, 46%, and 54% of the total per-step time for the 1.5B, 3B, and 7B models, respectively, with the remaining time primarily spent on policy updates [^3] In practice, for longer generation lengths, such as 8K and 16K, rollout time increases substantially, making it the dominant computational bottleneck. In such settings, our rollout replay mechanism can yield even greater wall-clock savings.. By reducing the number of rollouts per step, our RR strategy leads to a 11%–13% reduction in per-step training time, as shown in Tab. 1.
@@
As shown in Fig. 3, DOTS+RR (orange) consistently outperforms the original GRPO (blue) throughout training, maintaining higher accuracy at almost every step. Across all six settings, DOTS+RR reduces total training time by an average of 38.8%, with the largest improvement observed on Qwen2.5-3B trained on DeepMath (64.6%).
Table 1: Percentage of training steps saved, per-step time saved, and total training time saved. All metrics are averaged over four mathematical reasoning benchmarks and reported relative to the original GRPO baseline. All timing measurements are conducted on the same computational devices.
@@
Qwen2.5-Math-1.5B MATH 23.33 11.71 32.15
DeepScaleR 40.00 11.69 46.90
ORZ 20.00 11.66 29.20
@@
DeepMath 60.00 11.35 64.60
Qwen2.5-Math-7B DeepScaleR 13.33 13.39 24.94
@@ 5.3 Effectiveness of Adaptive Difficulty Prediction Framework
The adaptive difficulty prediction aligns with evolving training dynamics. To assess the fitness of online predictions, we collect ground-truth adaptive difficulty labels from training batches and compute the Pearson correlation between these labels and the predicted difficulty scores. As shown in Tab. 2, our framework consistently achieves strong Pearson correlation ($\rho>0.7$) across settings, demonstrating its ability to effectively track policy behavior throughout training. Additional qualitative examples are provided in Appendix D.2 for further insight into our attention-based prediction mechanism.
Figure 4: Ratio of effective questions (i.e., questions with adaptive difficulties strictly between 0 and 1) during training across various LLM-training dataset combinations. Annotated percentages indicate the per-step increase in effective question ratio achieved by DOTS compared to original GRPO, averaged over the training process. Our adaptive prediction framework consistently selects more informative samples throughout training.
Our prediction framework effectively filters out uninformative samples. Based on the discussion in §4.2, questions with adaptive difficulty 00 or $1$ correspond to cases where all rollouts receive the same reward. In such cases, the group-normalized advantage becomes zero, yielding no gradient signal. We define effective questions as those with adaptive difficulty strictly between 0 and 1. As shown in Fig. 4, on average across all LLM-dataset combinations, DOTS selects 25.4% more effective questions than the original GRPO, demonstrating a clear advantage in selecting more informative questions throughout training, thereby accelerating convergence.
@@ 5.4 Analysis and Discussion
We further investigate two important questions: (Q1) What are the individual contributions of DOTS and RR to training efficiency? (Q2) How does DOTS compare to a online data selection method based on external difficulty labels?
@@
DOTS accelerates convergence, while RR reduces per-step cost. As shown in Fig. 5(a), training guided by DOTS alone yields a steeper learning curve compared to the original GRPO algorithm. Fig. 5(b) shows that incorporating RR further reduces total training time by approximately 20% without sacrificing performance. These results show that DOTS and RR improve RL training efficiency in complementary ways.
@@
Figure 5: Average accuracy curves of (a) DOTS vs. Original GRPO, and (b) DOTS+RR vs. DOTS on the Qwen2.5-Math-1.5B model. The curves show average performance aggregated over four benchmarks with exponential smoothing. Note that the x-axis is the number of steps (rather than time). (a) DOTS consistently outperforms the original GRPO and leads to faster convergence. (b) Incorporating RR reduces training time by 20% while preserving the performance of DOTS.
@@
DOTS outperforms online data selection method based on external difficulty labels. We compare DOTS with an online data selection baseline that relies on external difficulty annotations (e.g., annotated by GPT-4o), where training questions are selected at different stages based on static difficulty labels, gradually shifting from easier to harder questions over time. Results in Appendix F.2 show that DOTS consistently outperforms this baseline. Moreover, such methods require expensive external labeling and offer limited adaptability, as they typically follow hand-crafted curricula that demand extensive manual design and tuning. In contrast, by leveraging adaptive difficulty, DOTS automatically adjusts to the model’s learning progress without relying on external supervision, enabling more scalable and efficient training.
@@ 6 Conclusion
In this paper, we propose two techniques to improve the data efficiency of LLM RL fine-tuning: difficulty-targeted online data selection and rollout replay. We hope these effective techniques encourage future work to explore data-centric approaches to improving LLM RL fine-tuning.
Appendix A Reproducibility
Our code repository is available at https://github.com/ASTRAL-Group/data-efficient-llm-rl/.
Appendix B Discussion
B.1 Limitations and Future Work
@@
Our adaptive difficulty prediction framework currently relies on randomly sampling a reference set of $K$ questions at each selection step. While effective, the quality of the reference set can influence prediction performance. In principle, one could improve prediction performance by selecting a more diverse reference set that better covers the dataset. Building on this idea, a natural extension is to fix a shared set of $K$ reference questions (with sufficient coverage) across training, re-evaluating their adaptive difficulty at each selection step.
@@
Moreover, while we demonstrate the effectiveness of experience replay in the GRPO setting, our current strategy is relatively straightforward: we randomly replay rollouts associated with questions whose average reward across all rollouts is neither 0 nor 1. A promising direction for further improving efficiency is to incorporate more principled strategies, such as drawing inspiration from prioritized experience replay [37, 51].
B.2 Extended Related Work
@@
RL fine-tuning of LLMs (with verifiable rewards) has recently attracted significant attention, driven in part by the success of DeepSeek-R1 [9]. Compared to the original GRPO algorithm [39], recent work has proposed several algorithmic improvements: DAPO [48] introduces techniques such as clip-higher, dynamic sampling, token-level policy gradient loss, and overlong reward shaping; Dr. GRPO [25] removes the length and standard deviation normalization terms to improve stability. Beyond these algorithmic enhancements, Vojnovic and Yun [42] and Mroueh [30] provide theoretical insights into GRPO, while Zeng et al. [50] and Yeo et al. [47] conduct large-scale empirical studies across models, sharing key design choices that enable RL fine-tuning.
@@
In contrast, relatively little attention has been paid to data-centric approaches. LIMR [22] explores a static data selection strategy for RL fine-tuning by prioritizing samples based on their alignment with the policy’s learning trajectory. However, it requires a full training run over the entire dataset beforehand, limiting its practicality. Our online data selection method DOTS is more efficient and applicable in realistic settings. In addition, prior work has not explored the use of rollout replay in GRPO, which we show can further reduce training costs.
Appendix C Proofs
@@
Consider a single question $q$, where $G$ responses $\{o_{i}\}_{i=1}^{G}$ are sampled independently from the current policy $\pi_{\theta}(\cdot\mid q)$. Each response receives a binary reward $r_{i}\in\{0,1\}$, sampled i.i.d. from a Bernoulli$(p)$ distribution, where $p$ represents the reward success rate. Define the group-relative advantage $\hat{A}_{i}$ as in Eq. 1. We consider the unclipped policy gradient estimator for this question without KL penalty:
@@
Assume that the likelihood gradients $\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)$ are independent of the reward distribution parameter $p$ and have bounded variance. Then, the expected squared norm of the gradient satisfies:
@@
Let $r_{i}\in\{0,1\}$ be the binary reward for response $o_{i}$, sampled i.i.d. from a Bernoulli$(p)$ distribution. Define the group-relative advantage as:
@@
Assume that the gradients $\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)$ are independent of the rewards $\{r_{i}\}$ and are independent across $i$, with bounded second moment:
@@
Because $\hat{A}_{i}$ and $\hat{A}_{j}$ are correlated, we compute the full second moment, where the expectation is taken with respect to $\pi_{\theta}$:
$\mathbb{E}[\|g\|^{2}]=\sum_{i,j=1}^{G}\mathbb{E}[\hat{A}_{i}\hat{A}_{j}]\cdot\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)^{\top}\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid q)].$
@@
Appendix D Details of Adaptive Difficulty Prediction Framework
D.1 Design and Implementation Details

### v2

@@ 5 Experiments
Figure 3: Average accuracy curves of our method and original GRPO under various LLM–dataset combinations. The curves show average performance aggregated over four benchmarks with exponential smoothing for visualization. The error bars represent 95% confidence intervals across 33 independent runs. Although both methods are trained for the same number of steps (60), our curve is shorter in duration because RR reduces the wall-clock time per step. Our method consistently outperforms the original GRPO throughout training and reduces the time required to match the original GRPO’s final accuracy after 60 training steps by an average of 40.7%.
@@ 5.1 Experimental Setup
We perform GRPO training on three model scales: Qwen2.5-Math-1.5B, Qwen2.5-3B, and Qwen2.5-Math-7B [46]. We adopt four open-source datasets for training: MATH [13], DeepScaleR-40K [27], Open-Reasoner-Zero-57K (ORZ) [15] and DeepMath-103K [12]. For MATH, we include all level 3–5 questions. For the other three datasets, we sample 8K to 10K subsets to construct the training pools. These datasets span diverse mathematical domains and difficulty levels. In total, we experiment with six LLM-training dataset combinations to assess the effectiveness of our framework.
@@
In practice, we observe that off-the-shelf pretrained embedding models struggle to capture fine-grained similarity between math questions. To address this, we freeze the Qwen2.5-Math-1.5B-Instruct backbone [46] and train a 3-layer MLP adapter with a calibration head, using binary cross-entropy loss. We fix the reference set size to 256. Additional implementation details are provided in Appendix C.1 and ablation results on key components are presented in Appendix E.1.
@@
We employ the verl framework [40] to perform GRPO training. We use a batch size of 512 and a mini-batch size of 64 in verl’s configuration, resulting in 8 gradient steps per training step. Across all experiments, we train for 60 training steps, yielding a total of 480 gradient steps. For each prompt, we generate 8 rollouts. The maximum rollout length is set to 3072 tokens for the Qwen2.5-Math series models (due to max position embedding limits) and 4096 tokens for Qwen2.5-3B. For reward computation, we use a simple rule-based function based solely on answer correctness, without incorporating any format-related signals. The 1.5B and 3B models are trained on 8 L40S GPUs, while the 7B model is trained on 8 A100 GPUs. For DOTS, data selection is performed every 2 steps. For RR, we choose the fresh rollout ratio $\delta$ as $0.5$ and buffer capacity $C\in\{256,512\}$. All RL training hyperparameters are detailed in Appendix D.2.
@@ 5.2 Main Results
The total training costs can be decomposed into two components: the number of steps required to reach a target performance and the average wall-clock time per step. Each training step involves processing a fixed-size batch, consisting primarily of rollout generation and policy update. To ensure a fair comparison, each set of experiments is run on the same type and number of GPU devices.
Our method reaches the same performance as the original GRPO with fewer steps.
Tab. 1 reports the number of training steps required by DOTS+RR to match the final performance of the original GRPO at 60 steps. Across all LLM–dataset combinations, our method consistently reaches the same performance with substantially fewer steps, achieving reductions ranging from 13.33% to 56.67%. These results demonstrate that DOTS significantly accelerates convergence by prioritizing informative training samples.
@@
In our experiments, rollout generation accounts for approximately 47%, 46%, and 54% of the total per-step time for the 1.5B, 3B, and 7B models, respectively, with the remaining time primarily spent on policy updates [^3] In practice, for longer generation lengths, such as 8K and 16K, rollout time increases substantially, making it the dominant computational bottleneck. In such settings, our rollout replay mechanism can yield even greater wall-clock savings.. By reducing the number of fresh rollouts per step, our RR strategy leads to a 11%–13% reduction in per-step training time, as shown in Tab. 1.
@@
As shown in Fig. 3, DOTS+RR (orange) consistently outperforms the original GRPO (blue) throughout training, maintaining higher accuracy at almost every step. Across all six settings, DOTS+RR reduces total training time by an average of 40.7%, with the largest improvement observed on Qwen2.5-3B trained on DeepMath (61.65%).
Table 1: Percentage of training steps saved, per-step time saved, and total training time saved. Results are averaged over four mathematical reasoning benchmarks and reported relative to the original GRPO baseline. All timing measurements are conducted on the same computational devices.
@@
Qwen2.5-Math-1.5B MATH 16.67 11.71 26.25
DeepScaleR 43.33 11.69 49.85
ORZ 13.33 11.66 23.30
@@
DeepMath 56.67 11.35 61.65
Qwen2.5-Math-7B DeepScaleR 40.00 13.39 48.03
@@ 5.3 Effectiveness of Adaptive Difficulty Prediction Framework
The adaptive difficulty prediction aligns with evolving training dynamics. To assess the fitness of online predictions, we collect ground-truth adaptive difficulty labels from training batches and compute the Pearson correlation between these labels and the predicted difficulty scores. As shown in Tab. 2, our framework consistently achieves strong Pearson correlation ($\rho>0.7$) across settings, demonstrating its ability to effectively track policy behavior throughout training. Additional qualitative examples are provided in Appendix C.2 to offer further insight into our attention-based prediction mechanism.
Our prediction framework effectively filters out uninformative samples. As discussed in §4.2, questions with adaptive difficulty values of 0 or $1$ correspond to cases where all rollouts receive identical reward. In such cases, the group-normalized advantage becomes zero, yielding no gradient signal. We define effective questions as those with adaptive difficulty strictly between 0 and 1. As shown in Fig. 4, on average across all LLM-dataset combinations, DOTS selects 25.4% more effective questions than the original GRPO, demonstrating a clear advantage in selecting more informative questions throughout training, thereby accelerating convergence.
@@ 5.4 Analysis and Discussion
We further investigate three important questions: (Q1) What are the individual contributions of DOTS and RR to training efficiency? (Q2) How does DOTS compare to an online data selection method based on external difficulty labels? (Q3) Do DOTS and RR remain effective in non-mathematical domains?
@@
DOTS accelerates convergence, while RR reduces per-step cost. As shown in Fig. 5(a), training guided by DOTS alone yields a steeper learning curve compared to original GRPO. Fig. 5(b) shows that incorporating RR further reduces training time by approximately 20% without sacrificing performance. These results show that DOTS and RR improve RL training efficiency in complementary ways.
Figure 4: Ratio of effective questions (i.e., questions with adaptive difficulties strictly between 0 and 1) during training across various LLM-training dataset combinations. Annotated percentages indicate the per-step increase in effective question ratio achieved by DOTS compared to original GRPO, averaged across the training process. Our adaptive prediction framework consistently selects more informative samples throughout training.
@@
Figure 5: Average accuracy curves of (a) DOTS vs. Original GRPO, and (b) DOTS+$R$ vs. DOTS on the Qwen2.5-Math-1.5B model. The curves show average performance aggregated over four benchmarks with exponential smoothing. Note that the x-axis is the number of steps (rather than time). (a) DOTS consistently outperforms the original GRPO and leads to faster convergence. (b) Incorporating RR reduces training time by 20% while preserving the performance of DOTS.
@@
DOTS outperforms online data selection method based on external difficulty labels. We compare DOTS with an online data selection baseline that relies on external difficulty annotations (e.g., annotated by GPT-4o-mini), where training questions are selected at different stages based on static difficulty labels, gradually shifting from easier to harder questions over time.
Specifically, we use the DeepScaleR dataset and label each question with GPT-4o-mini, following the difficulty annotation prompt introduced in [27]. Each question is annotated 32 times, and the average score is used as its final difficulty. We then follow a staged curriculum: in the first third of training steps, batches are sampled from the easiest third of the dataset; in the middle third, from the medium-difficulty third; and in the final third, from the hardest third. To ensure a fair comparison of online data selection strategies, we compare this baseline with DOTS (without RR). As shown in Fig. 6(a), our DOTS method consistently outperforms this baseline on both Qwen2.5-Math-1.5B and Qwen2.5-3B. Moreover, such methods require expensive external labeling and offer limited adaptability, as they typically follow hand-crafted curricula that demand extensive manual design and tuning. In contrast, by leveraging adaptive difficulty, DOTS automatically adjusts to the model’s learning progress without relying on external supervision, enabling more scalable and efficient training.
(a) Comparison between DOTS (ours) and an external difficulty-based curriculum baseline. The curves show average performance aggregated over four benchmarks with exponential smoothing for visualization. Note that the x-axis is the number of steps (rather than time). Our method consistently outperforms the baseline.
(b) Results on the science subsets of MMLU using Qwen2.5-3B trained on the SCP-25K dataset. Our method significantly outperforms original GRPO in this non-math domain.
Figure 6: Comparison of DOTS with external difficulty-based curriculum baseline (left) and generalization to non-math domain (right).
DOTS and RR improve RL data efficiency beyond mathematics.
To further examine the generality of our approach beyond the math domain, we apply the full training and evaluation pipeline to the science domain using the curated SCP-25K dataset [liu2025prorl], which mostly contains advanced physics, chemistry, and biology questions. We adopt the Qwen2.5-3B model and train a new adaptive difficulty predictor, while keeping all other RL settings unchanged. We evaluate performance on the science subsets of MMLU, including questions in the fields of physics, chemistry, and biology. As reported in Fig. 6(b), our method continues to significantly improve RL data efficiency in this non-math domain, demonstrating its broader applicability.
@@ 6 Conclusion
In this paper, we propose two techniques to improve the data efficiency of LLM RL fine-tuning: Difficulty-targeted Online Data Selection and Rollout Replay. We hope these effective techniques will encourage future work to explore data-centric approaches to improving LLM RL fine-tuning.
Acknowledgment
We would like to express our heartfelt thanks to Rayne Amami for helpful discussions and inputs.
Appendix A Discussions
A.1 Limitations and Future Work
@@
Our adaptive difficulty prediction framework currently relies on randomly sampling a reference set of $K$ questions at each selection step. While effective, the quality of the reference set can influence prediction performance. In principle, one could improve prediction performance by selecting a more diverse reference set that better covers the training set. Building on this idea, a natural extension is to fix a shared set of $K$ reference questions (with sufficient coverage) across training, re-evaluating their adaptive difficulty at each selection step.
@@
Moreover, while we demonstrate the effectiveness of experience replay in the GRPO setting, our current strategy is relatively straightforward: we randomly replay rollouts associated with questions whose average reward across all rollouts is neither 0 nor 1. A promising direction for further improving efficiency is to incorporate more principled replay strategies, such as those inspired prioritized experience replay [37, 51].
Another potential extension of our method lies in the construction of input embeddings for difficulty prediction. Specifically, instead of relying solely on the question text, one could incorporate reference solutions to enrich the representation. Preliminary experiments suggest that including reference solutions can slightly improve the accuracy of adaptive difficulty prediction. However, this approach may have limited applicability in practice, as reference solutions are not available for all datasets (e.g., DeepScaler and ORZ).
Finally, we note that generating rollouts for the reference set can introduce nontrivial computational overhead, especially when the reference size is large. To mitigate this, we reuse rollouts from reference questions whose predicted difficulty is near 0.5, effectively incorporating them into training. This strategy reduces rollout generation cost by 4–12% per step while maintaining final performance.
A.2 Extended Related Work
@@
RL fine-tuning for LLMs (with verifiable rewards) has recently attracted significant attention, driven in part by the success of DeepSeek-R1 [9]. Compared to the original GRPO algorithm [39], recent work has proposed several algorithmic improvements: DAPO [48] introduces techniques such as clip-higher, dynamic sampling, token-level policy gradient loss, and overlong reward shaping, while Dr. GRPO [25] removes the length and standard deviation normalization terms to improve stability. Beyond these algorithmic enhancements, [42, 30] provide theoretical insights into GRPO, while [50, 47] conduct large-scale empirical studies across models, identifying key design choices that enable effective RL fine-tuning.
@@
In contrast, relatively little attention has been paid to data-centric approaches, despite their demonstrated potential in other areas of LLM training [kang2024get, kang2024autoscale, deng2025survey]. LIMR [22] explores a static data selection strategy for RL fine-tuning by prioritizing samples based on their alignment with the policy’s learning trajectory. However, it requires a full training run over the entire dataset beforehand, limiting its practicality. Our online data selection method DOTS is more efficient and applicable in realistic settings. In addition, prior work has not explored the use of rollout replay in LLM RL fine-tuning, which we show can further reduce training costs.
Appendix B Proofs
@@
Consider a single question $q$, where $G$ responses $\{o_{i}\}_{i=1}^{G}$ are sampled independently from the current policy $\pi_{\theta}(\cdot\mid q)$. Each response receives a binary reward $r_{i}\in\{0,1\}$, sampled i.i.d. from a Bernoulli(p)(p) distribution, where $p$ represents the reward success rate. Define the group-relative advantage $\hat{A}_{i}$ as in Eq. 1. We consider the unclipped policy gradient estimator for this question without KL penalty:
@@
Under mild assumptions on the reward and the likelihood gradients $\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)$, the expected squared norm of the gradient satisfies:
@@
Let $r_{i}\in\{0,1\}$ be the binary reward for response $o_{i}$, sampled i.i.d. from a Bernoulli(p)(p) distribution. Define the group-relative advantage as:
@@
Assume that the gradients $\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)$ have bounded second moment:
@@
We compute the full second moment of the gradient estimator, where the expectation is taken with respect to $\pi_{\theta}$:
$\mathbb{E}[\|g\|^{2}]$ $=\underbrace{\sum_{i,j=1}^{G}\mathbb{E}\left[\hat{A}_{i}\hat{A}_{j}\right]\cdot\mathbb{E}\left[\nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)^{\top}\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid q)\right]}_{T_{1}}$
$\qquad+\underbrace{\sum_{i,j=1}^{G}\mathrm{Cov}\left(\hat{A}_{i}\hat{A}_{j},\ \nabla_{\theta}\log\pi_{\theta}(o_{i}\mid q)^{\top}\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid q)\right)}_{T_{2}}.$
We introduce a weak-dependence assumption that the correction term $T_{2}$ is negligible compared to the leading term $T_{1}$ [^4] To support this assumption empirically, we compute the ratio $\left|\frac{T_{2}}{T_{1}}\right|$ on two LLM-dataset combinations. Specifically, we randomly sample 512 questions for each dataset, generate 8 rollouts per question, and evaluate the ratio. The results (mean $\pm$ standard deviation) are: Qwen2.5-Math-1.5B + MATH: $0.081\pm 0.0065$, and Qwen2.5-3B + DeepMath: $0.081\pm 0.0051$. These consistently low ratios empirically validate the weak-dependence assumption.:
$\left|\frac{T_{2}}{T_{1}}\right|\ll 1.$
Therefore, it suffices to focus our analysis on the leading term $T_{1}$.
@@
Remark: Extension to Multi-component Rewards.
Theorem 1 focuses on binary rewards for simplicity, following standard practice in recent RLVR literature. Its core derivation—computing the second central moment of group-normalized rewards $\mathbb{E}[\hat{A}_{i}^{2}]$—extends naturally to more complex reward formulations.
For example, consider a reward composed of two independent components: a correctness term $c_{i}\sim\text{Bern}(\alpha)$ and a format term $f_{i}\sim\text{Bern}(\beta)$, where the total reward is $r_{i}=c_{i}+f_{i}$. Then,
$\mathbb{E}[\hat{A}_{i}^{2}]=\left(\alpha(1-\alpha)+\beta(1-\beta)\right)\left(1-\frac{1}{G}\right),$
which is maximized when both $\alpha=0.5$ and $\beta=0.5$. This demonstrates that our insight applies naturally to multi-component rewards and highlights the generality of the result.
Appendix C Details of Adaptive Difficulty Prediction Framework
C.1 Design and Implementation Details